OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

y +2 = 2/5 (x -5)

OpenStudy (anonymous):

get that in the format of y = mx + b by distribution and isolating Y

OpenStudy (anonymous):

they want it in a form like this _ x _ _ y _ _ _

OpenStudy (anonymous):

ok - just isolate the variables x and y on one side - you still must distribute

OpenStudy (anonymous):

lol how do i do that

hero (hero):

by the way, the general formula for that is: y -y1 = m(x - x1) Substitute the slope, m = 2/5, and the points (x1,y1) = (5,-2) into the formula, then solve for y. Don't forget to simplify properly.

OpenStudy (anonymous):